Can you show me how to find an equation of the line containing the given pair of points (4,3) and (12,8)?
Answer by stanbon(75887)

find an equation of the line containing the given pair of points (4,3) and (12,8)?
------
slope = (3-8)/(4-12) = -5/-8 = 5/8
----
Form: y = mx + b
----
Solve for "b":
3 = (5/8)4 + b
3 = 5/2 + b
b = 1/2
-------------------
Equation:
y = (5/8)x + (1/2)
=====================